# On the dimension distortion under fractionally smooth mappings

### v1

@@ 4. Intermediate dimensions under fractionally smooth Sobolev mappings
there is $C_{E}=C_{K}^{t}C_{K}^{\prime}\tilde{C}\left(\int_{K}\tilde{g}\,d\mu\right)<\infty$ such that
@@
Suppose that $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}$ is finite and let $\vec{g}:=\{g_{k}\}_{k\in\mathbb{Z}}\in\mathbb{D}^{s}(f)\cap\ell^{q}(L^{p}(X))$. By Lemma 4.5(3), if $q\in(0,p]$, then $\dot{N}^{s}_{p,q}(X:Y)\hookrightarrow\dot{M}^{s,p}(X:Y)$ and the desired conclusions follow from what we have already proven. Thus, suppose $q>p$ and fix $t\in(\min\{1,Q/s\},p)$. Then it follows from Hölder’s inequality that $\|f\|_{\dot{N}^{s}_{t,q}(B:Y)}$ is finite for each fixed ball $B\subset X$, where the pointwise restriction of the functions in the sequence $\{g_{k}\}_{k\in\mathbb{Z}}$ to $B$ serves as a fractional $s$-gradient of $f$. By this, (4.28), and Corollary 4.6, we have that there exists a constant $C_{k}\in[1,\infty)$, which is independent of $f$, such that, for all balls $B_{i}:=B(x_{i},r)$, the following inequality holds
@@
for some sufficiently large measurable set $K^{\prime}\subset X$ containing $K$ and some constant $C^{\prime}\in(0,\infty)$ that only depends on the doubling constant $\tilde{C}_{\text{hom}}$ from (2.14). This, combined with (4.2) and the fact that (keeping in mind $\vec{g}:=\{g_{k}\}_{k\in\mathbb{Z}}\in\ell^{q}(L^{p}(X))$)
$\sum_{k\in\mathbb{Z}}\left(\int_{K}g_{k}^{p}\,d\mu\right)^{q/p}=\|\vec{g}\|_{\ell^{q}(L^{p}(K))}^{q}\leq\|\vec{g}\|_{\ell^{q}(L^{p}(X))}^{q}<\infty,$
@@
$\sum_{i\in I}\sum_{k\in\mathbb{Z}}\left(\int_{\varepsilon B_{i}}\tilde{g}_{k}\,d\mu\right)^{q/t}\leq(C^{\prime})^{q/t}[\mu(K)]^{q(1-t/p)/t}\sum_{k\in\mathbb{Z}}\left(\int_{K}g_{k}^{p}\,d\mu\right)^{q/p}<\infty.$
Thus, for $C_{E}:=C_{K}^{q}(C_{K}^{\prime}\tilde{C}C^{\prime})^{q/t}[\mu(K)]^{q(1-t/p)/t}\sum_{k\in\mathbb{Z}}\left(\int_{K}g_{k}^{p}\,d\mu\right)^{q/p}<\infty$, we have
@@ 5. Hausdorff and Minkowski dimension distortion
We first note that $X$ supporting a $Q$-Poincaré inequality implies that $X$ is connected [47, Proposition 8.1.6] and, thus, uniformly perfect. As a result, Corollary 1.2 (i), (ii) are direct implications of [21, Theorem 1.2] and [46, Theorem 9.3], respectively. See also the proof of [21, Corollary 1.3] for more details.
@@
(ii): This follows directly by [21, Theorem 1.1] and Theorem 1.3. ∎
@@
It should be noted that in the case where $X,Y$ are uniformly perfect, letting $\theta\rightarrow 1$ in (1.1) yields Theorem 1.4 (ii), without the use of [21, Theorem 1.1]. For Theorem 1.4 (i), although the proof is a very simplified case of that of Theorem 1.1 for $\theta=0$, simply letting $\theta\rightarrow 0$ in (1.1) does not generally yield the desired inequality. This is due to the intermediate dimensions not necessarily being continuous at $\theta=0$ (see for instance [28]).

### v2

@@ 4. Intermediate dimensions under fractionally smooth Sobolev mappings
there is a constant $C_{E}:=C_{K}^{t}C_{K}^{\prime}\tilde{C}\left(\int_{K}\tilde{g}\,d\mu\right)<\infty$ satisfying
@@
Suppose that $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}$ is finite. Note that $\|f\|_{\dot{N}^{s}_{p,q}(K^{\prime}:Y)}$ is finite and so there exists $\vec{g}:=\{g_{k}\}_{k\in\mathbb{Z}}\in\mathbb{D}^{s}(f)\cap\ell^{q}(L^{p}(K^{\prime}))$. By Lemma 4.5(3), if $q\in(0,p]$, then $\dot{N}^{s}_{p,q}(K^{\prime}:Y)\hookrightarrow\dot{M}^{s,p}(K^{\prime}:Y)$ and the desired conclusions follow from what we have already proven. Thus, suppose $q>p$ and fix $t\in(Q/s,p)$. Similar to as before, we have that $\|f\|_{\dot{N}^{s}_{t,q}(2B_{i}:Y)}$ is finite for each fixed $B_{i}$, where the pointwise restriction of the functions in the sequence $\{g_{k}\}_{k\in\mathbb{Z}}$ to $2B_{i}$ serves as a fractional $s$-gradient of $f$. Again, we view each $g_{k}:K^{\prime}\to[0,\infty]$ as a locally $p$-integrable function defined on $X$ by extending it by zero outside its original domain. By this, (4.28), and Corollary 4.6, we have that there exists a constant $C_{K}\in[1,\infty)$, which is independent of $f$, such that, for all balls $B_{i}$, the following inequality holds
@@
for some constant $C^{\prime}\in(0,\infty)$ that only depends on the doubling constant $\tilde{C}_{\text{hom}}$ from (2.13). This, combined with (4.2) and the fact that
$\sum_{k\in\mathbb{Z}}\left(\int_{K^{\prime}}g_{k}^{p}\,d\mu\right)^{q/p}=\|\vec{g}\|_{\ell^{q}(L^{p}(K^{\prime}))}^{q}<\infty,$
@@
$\sum_{i\in I}\sum_{k\in\mathbb{Z}}\left(\int_{\varepsilon B_{i}}\tilde{g}_{k}\,d\mu\right)^{q/t}\leq(C^{\prime})^{q/t}[\mu(K)]^{q(1-t/p)/t}\sum_{k\in\mathbb{Z}}\left(\int_{K^{\prime}}g_{k}^{p}\,d\mu\right)^{q/p}<\infty.$
Thus, for $C_{E}:=C_{K}^{q}(C_{K}^{\prime}\tilde{C}C^{\prime})^{q/t}[\mu(K)]^{q(1-t/p)/t}\sum_{k\in\mathbb{Z}}\left(\int_{K^{\prime}}g_{k}^{p}\,d\mu\right)^{q/p}<\infty$, we have
@@
Remark 4.9.
A closer study of the proof of Theorem 1.3 above reveals that essentially the finiteness of the respective semi-norms was only utilized when restricted to the compact set $K$. Thus, both the inclusion in the compactly Hölder class and the dimension bounds are in fact true even for continuous mappings in the local Hajłasz–Triebel–Lizorkin and the local Hajłasz–Besov classes.
@@ 5. Hausdorff and Minkowski dimension distortion
We first note that $X$ supporting a $Q$-Poincaré inequality implies that $X$ is connected [49, Proposition 8.1.6] and, thus, uniformly perfect. As a result, Corollary 1.2 (i), (ii) are direct implications of [22, Theorem 1.2] and [48, Theorem 9.3], respectively (see Remark 2.5). See also the proof of [22, Corollary 1.3] for more details.
@@
(ii): This follows directly by [22, Theorem 1.1] and Theorem 1.3. ∎
@@
It should be noted that in the case where $X,Y$ are uniformly perfect, letting $\theta\rightarrow 1$ in (1.1) yields Theorem 1.4 (ii), without the use of [22, Theorem 1.1]. For Theorem 1.4 (i), although the proof is a very simplified case of that of Theorem 1.1 for $\theta=0$, simply letting $\theta\rightarrow 0$ in (1.1) does not generally yield the desired inequality. This is due to the intermediate dimensions not necessarily being continuous at $\theta=0$ (see for instance [30]).